OpenStudy (anonymous):

In the solution of the equation 5 - 3x = 2x + 9, 3x is added to the equation first. Which of the following should be done next? add -9 add -5 add -2x

OpenStudy (anonymous):

5 - 3x = 2x + 9 + 3x +3x 5 = 5x + 9

OpenStudy (anonymous):

so add -9 is the answer

OpenStudy (anonymous):

I agree. Add -9.

OpenStudy (ybarrap):

$$ \large{ 5 - 3x = 2x + 9\\ 1)~ 5 - 3x+3x = 2x + 9+3x\\ 2)~5-9=5x+9-9\\ 3)~-4=5x\\ 4)x={-4\over 5} } $$ Does this make sense?